OpenStudy (anonymous):

If x+1/x=-5, what is x^3+1/x^3?

OpenStudy (anonymous):

I will give a medal.

Parth (parthkohli):

\[\left(x + \dfrac{1}{x}\right)^3 = x^3 + \dfrac{1}{x^3} + 3\left(x + \dfrac{1}{x}\right)\]

OpenStudy (anonymous):

so what do i do after that?

Parth (parthkohli):

You use it.

OpenStudy (anonymous):

but how?

OpenStudy (anonymous):

I don't really understand how to even approach this problems. Can you please walk me through it?

OpenStudy (anonymous):

@ParthKohli ? are you still there?